# On the Value of Cross-Modal Misalignment in Multimodal Representation Learning

### v4

@@ B.1 Lemmas
On the other hand, by the pairing conditions in Asm. 3.2, there exists at least one subset $A\subseteq\mathbb{I}_{\rho}$ of perturbed semantic variables such that $l\in A$ and $p_{A}(A)>0$. Pick one such set and call it $A$. Define the latent semantic variables corresponding to the image of this pair as $(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$ (here, we omit the latent semantic components that have already been excluded in previous steps).
By Asm. 3.2, we know that $\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}=\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}\setminus A}$ a.e., that is, $({\mathbf{s}}_{\mathbb{I}_{\rho}\setminus A},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\rho}})=(\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}\setminus A},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})$ a.e.. Thus, we can rewrite text semantic variables $\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}}=(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})$ as $(\tilde{\mathbf{s}}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$.
Further, also by Asm. 3.2, there exists a non-empty open subspace $\mathcal{O}_{A}\subseteq\mathcal{S}_{A}$ such that $p_{\tilde{\mathbf{s}}_{A}|\mathbf{s}_{A}}(\cdot|\mathbf{s}^{*}_{A})$ is strictly positive on $\mathcal{O}_{A}$. Since the perturbed random variable $\tilde{\mathbf{s}}^{*}_{A}$ is a realization within this open subspace, we know it lies in $\mathcal{O}_{A}$ and $\mathcal{O}_{A}$ is non-empty. Moreover, because $p_{\tilde{\mathbf{s}}_{A}|\mathbf{s}_{A}}(\cdot|\mathbf{s}^{*}_{A})$ is smooth and strictly positive on $\mathcal{O}_{A}$, there exists a sufficiently small $\eta_{1}>0$ such that
@@
Based on the invariance condition established in Eq. 24, we have the following equalities:
@@
This is contradicted by the inequality established in Eq. 25, which implies that such a $\mathbf{h}^{c}_{t}$ cannot exist. Consequently, any $\mathbf{h}_{t}$ minimizing the loss must be independent of the perturbed semantic variables $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$, i.e.,
@@
Updating the invariance condition, and combining Equations 24 and 26, we obtain:
@@
The exclusion of $\mathbf{s}_{\mathbb{I}_{\rho}}$ from image representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$ can be demonstrated by a similar procedure to Step 3, namely, that excluded semantic components from one modality cannot exist in another view, regardless of the latent causal structure among $\mathbf{s}_{\mathbb{I}_{\theta}}$. Specifically, fixing the value of $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ and varying $\mathbf{s}_{\mathbb{I}_{\rho}}$ within a small region, we can sample distinct semantic variables (due to $p_{\mathbf{s}}>0$ a.e. over $\mathcal{S}$ by Asm. 3.1). The smoothness of $\mathbf{h}_{x}$ then leads to an inequality if it depends on any component in $\mathbb{I}_{\rho}$. This inequality contradicts the alignment condition established in Eq. 27. Thus, $\mathbf{h}_{x}$ must also be independent of $\mathbf{s}_{\mathbb{I}_{\rho}}$.
@@
(Preservation of All Unbiased Semantic Variables). Based on Equations 28 and 29, we define the learned image and textual representations as
@@
According to Equations 10 and 11, we know that the learned representations in both modalities are uniformly distributed over $(0,1)^{n}$. By directly applying Lem. B.2, it follows that the learned representations $\hat{\mathbf{z}}_{x}$ (and also $\hat{\mathbf{z}}_{t}$) include all and only the information of the unaltered semantic components $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ a.s., and that $\mathbf{h}_{x}$ (and similarly $\mathbf{h}_{t}$) is invertible. Consequently, the true modality-shared semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ are block-identified by $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
Thus, the proof of Thm. 4.1 is complete. ∎
B.3 Proof of Corollary 4.1
We now proceed to prove Cor. 4.1. To begin, we restate the corollary for clarity: See 4.1
@@
As we have fixed a graded lexicographic order over the range of $\theta$, that is, over $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ as defined in Defn. 3.1. Then, setting $\theta=2^{n_{s}}-1$ corresponds to selecting the full set of semantic variables for text generation, i.e., $\mathbb{I}_{\theta}=\mathbb{I}_{\mathbf{s}}$.
Furthermore, we have similarly fixed a graded lexicographic order over the range of $\rho$, i.e., over $\mathcal{P}_{\text{proper}}(\mathbb{I}_{\mathbf{s}})$, as defined in Defn. 3.2. Given that $\mathbb{I}_{\theta}=\mathbb{I}_{\mathbf{s}}$, setting $\rho=1$ implies that all semantic variables $\mathbf{s}$ are preserved without perturbation during the generation of the corresponding text $\mathbf{t}^{\scriptscriptstyle(\theta)}$.
Under these assumptions, and by Asm. 3.2, the perturbing subset $A$ is always trivial because $\mathbb{I}_{\rho}$ is trivial. Consequently, we have
@@
Now consider any pair of smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $\mathbf{f}_{t^{\scriptscriptstyle(\theta)}}:\mathcal{T^{\scriptscriptstyle(\theta)}}\to(0,1)^{n_{s}}$ that achieve the global minimum of the loss in Eq. 2, i.e., yield a value of zero. From Step 1 and Step 2 of the proof of Thm. 4.1, it follows that:
@@
Since both the omitted and perturbable index subsets are trivial, we may directly apply Lem. B.2, which implies that the full semantic vector $\mathbf{s}$ is block-identified by both $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
@@
Clarification B.2 (Fixed Graded Lexicographic Order).
The graded lexicographic order over subsets of semantic indices is fixed solely for notational consistency and clarity. For example, given semantic variables $\mathbf{s}=(s_{1},s_{2},s_{3})$ with index set $\mathbb{I}_{\mathbf{s}}=\{1,2,3\}$, the graded lexicographic order over the non-empty powerset $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})$ yields: $\mathcal{P}_{+}(\mathbb{I}_{\mathbf{s}})=\{\{1\},\{2\},\{3\},\{1,2\},\{1,3\},\{2,3\},\{1,2,3\}\}.$ This ordering imposes no constraints on the underlying latent structure and is adopted without loss of generality. Since the true permutation of latent variables is unidentifiable, any consistent order can be used to index subset-valued parameters such as $\theta$ (selection) and $\rho$ (perturbation). Importantly, both $\theta$ and $\rho$ act indirectly through text observations rather than directly operating on the latent space. The fixed order simply determines how each value of $\theta$ or $\rho$ maps to a subset of semantic indices. For instance, omitting color in a caption corresponds to removing the associated latent variable—which index encodes this depends on the fixed ordering. Thus, adopting a graded lexicographic order ensures a deterministic and reproducible indexing scheme without limiting model generality or expressiveness.
B.4 Proof of Corollary 4.2
We now proceed to prove Cor. 4.2. For clarity, we restate the corollary below: See 4.2
@@
Given this structure, we directly apply the argument from the proof of Thm. 4.1. The omission of variables in $\mathbb{I}^{2}_{\text{var}}$, together with the perturbation of variables in $\mathbb{I}^{1}_{\text{var}}$, ensures that only the invariant subset $\mathbb{I}_{\text{inv}}$ is both selected and unperturbed. Therefore, the invariant semantic components $\mathbf{s}_{\mathbb{I}_{\text{inv}}}$ are block-identified via smooth functions
@@
Appendix C Numerical Simulation Details
We provide additional details on the numerical simulations that are not fully covered in § 5.1. Specifically:
  • In § C.1, we outline the experimental setup, including hyperparameters, model architectures, and the construction of downstream tasks.
  • In § C.2, we present additional experiments that further validate our theoretical findings.
  • In § C.3, we analyze downstream performance under various perturbation bias settings.
C.1 Detailed Experimental Setup
@@
Given the 10-dimensional semantic space, the total number of nonempty subsets is $10232^{10}-1=$ 1023, defining the full range of selection bias configurations $\theta$. For tractability, we choose 10 representative selection subsets $\mathbb{I}_{\theta}$, detailed in Table 2.
To isolate perturbation bias, we fix full selection ($\theta=1023$) and evaluate 10 representative perturbation subsets $\mathbb{I}_{\rho}$, as listed in Table 3. Perturbations are introduced by independently modifying each semantic variable $s_{i}$ in $\mathbb{I}_{\rho}$ with probability 0.75 using additive Gaussian noise:
@@
where the joint noise covariance $\Sigma_{\epsilon}$ is sampled from $\mathcal{W}_{10}(I,10)$. Perturbations are applied solely to the text modality, aligning the proposed LVM and assumptions stated in § 3.
@@
$\rho\mid\theta=1023$ $1$ $2$ $12$ $57$ $177$ $387$ $639$ $849$ $969$ $1014$
$\mathbb{I}_{\rho}$emptyset $\{1\}$ $[2]$ $[3]$ $[4]$ $[5]$ $[6]$ $[7]$ $[8]$ $[9]$
@@
We parameterize the generative functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ using randomly initialized 3-layer invertible MLPs, following prior work [12]. Invertibility is enforced by maintaining a condition number threshold of $31\mathrm{e}{-}$ 3 for each layer.
The encoding functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ are implemented as 7-layer MLPs and optimized using the Adam optimizer. For MMCL training, we use a batch size of $6144$, a learning rate of $41\mathrm{e}{-}$ 4, and train for $100{,}000$ iterations. The loss function is given by Eq. 2, with Euclidean distance used as the similarity metric and a temperature parameter set to $1.0$. To ensure training stability, gradients are clipped using a maximum 2-norm of $2$.
@@
For both downstream tasks, we fix the pretrained encoders and evaluate the quality of the learned representations using a two-layer MLP as a probing model. We generate $20{,}480$ samples as the evaluation set for training the regressors and classifiers, along with an additional $20{,}480$ samples as the in-distribution test set. To assess OOD generalization, we generate another $20{,}480$ samples from the shifted latent space as the OOD test set.
The regressors are trained using Mean Squared Error (MSE) loss, and the classifiers use Cross-Entropy loss, both with a learning rate of $10^{-3}$ and trained for $10{,}000$ steps. We report classification performance using the average Matthews Correlation Coefficient (MCC). Importantly, in the OOD setting, we perform no adaptation or fine-tuning; the classifier is evaluated directly to test the generalization capability of the pretrained representations.
C.2 Additional Identification Results

### v5

@@ C.1 Lemmas
On the other hand, by the pairing conditions in Asm.˜3.2, there exists at least one subset $A\subseteq\mathbb{I}_{\rho}$ of perturbed semantic variables such that $l\in A$ and $p_{A}(A)>0$. Pick one such set and call it $A$. Define the latent semantic variables corresponding to the image of this pair as $(\mathbf{s}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$ (here, we omit the latent semantic components that have already been excluded in previous steps).
By Asm.˜3.2, we know that $\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A}=\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}\setminus A}$ a.e., that is, $({\mathbf{s}}_{\mathbb{I}_{\rho}\setminus A},\mathbf{s}^{*}_{\mathbb{I}^{c}_{\rho}})=(\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}\setminus A},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})$ a.e.. Thus, we can rewrite text semantic variables $\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\theta}}=(\tilde{\mathbf{s}}^{*}_{\mathbb{I}_{\rho}},\tilde{\mathbf{s}}^{*}_{\mathbb{I}^{c}_{\rho}})$ as $(\tilde{\mathbf{s}}^{*}_{A},\mathbf{s}^{*}_{\mathbb{I}_{\theta}\setminus A})$.
Further, also by Asm.˜3.2, there exists a non-empty open subspace $\mathcal{O}_{A}\subseteq\mathcal{S}_{A}$ such that $p_{\tilde{\mathbf{s}}_{A}|\mathbf{s}_{A}}(\cdot|\mathbf{s}^{*}_{A})$ is strictly positive on $\mathcal{O}_{A}$. Since the perturbed random variable $\tilde{\mathbf{s}}^{*}_{A}$ is a realization within this open subspace, we know it lies in $\mathcal{O}_{A}$ and $\mathcal{O}_{A}$ is non-empty. Moreover, because $p_{\tilde{\mathbf{s}}_{A}|\mathbf{s}_{A}}(\cdot|\mathbf{s}^{*}_{A})$ is smooth and strictly positive on $\mathcal{O}_{A}$, there exists a sufficiently small $\eta_{1}>0$ such that
@@
Based on the invariance condition established in Eq.˜24, we have the following equalities:
@@
This is contradicted by the inequality established in Eq.˜25, which implies that such a $\mathbf{h}^{c}_{t}$ cannot exist. Consequently, any $\mathbf{h}_{t}$ minimizing the loss must be independent of the perturbed semantic variables $\tilde{\mathbf{s}}_{\mathbb{I}_{\rho}}$, i.e.,
@@
Updating the invariance condition, and combining Equations˜24 and 26, we obtain:
@@
The exclusion of $\mathbf{s}_{\mathbb{I}_{\rho}}$ from image representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$ can be demonstrated by a similar procedure to Step 3, namely, that excluded semantic components from one modality cannot exist in another view, regardless of the latent causal structure among $\mathbf{s}_{\mathbb{I}_{\theta}}$. Specifically, fixing the value of $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ and varying $\mathbf{s}_{\mathbb{I}_{\rho}}$ within a small region, we can sample distinct semantic variables (due to $p_{\mathbf{s}}>0$ a.e. over $\mathcal{S}$ by Asm.˜3.1). The smoothness of $\mathbf{h}_{x}$ then leads to an inequality if it depends on any component in $\mathbb{I}_{\rho}$. This inequality contradicts the alignment condition established in Eq.˜27. Thus, $\mathbf{h}_{x}$ must also be independent of $\mathbf{s}_{\mathbb{I}_{\rho}}$.
@@
(Preservation of All Unbiased Semantic Variables). Based on Equations˜28 and 29, we define the learned image and textual representations as
@@
According to Equations˜10 and 11, we know that the learned representations in both modalities are uniformly distributed over $(0,1)^{n}$. By directly applying Lem.˜C.2, it follows that the learned representations $\hat{\mathbf{z}}_{x}$ (and also $\hat{\mathbf{z}}_{t}$) include all and only the information of the unaltered semantic components $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ a.s., and that $\mathbf{h}_{x}$ (and similarly $\mathbf{h}_{t}$) is invertible. Consequently, the true modality-shared semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ are block-identified by $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
Thus, the proof of Thm.˜4.1 is complete. ∎
C.3 Proof of Corollary 4.1
We now proceed to prove Cor.˜4.1. To begin, we restate the corollary for clarity: See 4.1
@@
As we have fixed a graded lexicographic order over the range of $\theta$, that is, over $\mathcal{P}^{+}(\mathbb{I}_{\mathbf{s}})$ as defined in Defn.˜3.1. Then, setting $\theta=2^{n_{s}}-1$ corresponds to selecting the full set of semantic variables for text generation, i.e., $\mathbb{I}_{\theta}=\mathbb{I}_{\mathbf{s}}$.
Furthermore, we have similarly fixed a graded lexicographic order over the range of $\rho$, i.e., over $\mathcal{P}_{\text{prop}}(\mathbb{I}_{\mathbf{s}})$, as defined in Defn.˜3.2. Given that $\mathbb{I}_{\theta}=\mathbb{I}_{\mathbf{s}}$, setting $\rho=1$ implies that all semantic variables $\mathbf{s}$ are preserved without perturbation during the generation of the corresponding text $\mathbf{t}^{\scriptscriptstyle(\theta)}$.
Under these assumptions, and by Asm.˜3.2, the perturbing subset $A$ is always trivial because $\mathbb{I}_{\rho}$ is trivial. Consequently, we have
@@
Now consider any pair of smooth functions $\mathbf{f}_{x}:\mathcal{X}\to(0,1)^{n_{s}}$ and $\mathbf{f}_{t^{\scriptscriptstyle(\theta)}}:\mathcal{T^{\scriptscriptstyle(\theta)}}\to(0,1)^{n_{s}}$ that achieve the global minimum of the loss in Eq.˜2, i.e., yield a value of zero. From Step 1 and Step 2 of the proof of Thm.˜4.1, it follows that:
@@
Since both the omitted and perturbable index subsets are trivial, we may directly apply Lem.˜C.2, which implies that the full semantic vector $\mathbf{s}$ is block-identified by both $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
@@
Definition C.1 (Graded Lexicographic Order).
Let $S$ be a finite set with a total order (e.g., $S=\{1,2,\dots,n\}$). The graded lexicographic order on the non-empty subsets of $S$ is defined as follows: (i) Subsets are ordered first by increasing cardinality (i.e., all subsets of size $k$ precede those of size $k+1$). (ii) Within each group of equal cardinality, subsets are ordered lexicographically: that is, $\{i_{1},i_{2},\dots,i_{k}\}<\{j_{1},j_{2},\dots,j_{k}\}$ if there exists an $\ell$ such that $i_{m}=j_{m}$ for all $m<\ell$ and $i_{\ell}<j_{\ell}$.
Clarification C.2 (Fixed Graded Lexicographic Order).
The graded lexicographic order over subsets of semantic indices is fixed solely for notational consistency and clarity. For example, given semantic variables $\mathbf{s}=(s_{1},s_{2},s_{3})$ with index set $\mathbb{I}_{\mathbf{s}}=\{1,2,3\}$, the graded lexicographic order over the non-empty powerset $\mathcal{P}^{+}(\mathbb{I}_{\mathbf{s}})$ yields: $\mathcal{P}^{+}(\mathbb{I}_{\mathbf{s}})=\{\{1\},\{2\},\{3\},\{1,2\},\{1,3\},\{2,3\},\{1,2,3\}\}.$ This ordering imposes no constraints on the underlying latent structure and is adopted without loss of generality. Since the true permutation of latent variables is unidentifiable, any consistent order can be used to index subset-valued parameters such as $\theta$ (selection) and $\rho$ (perturbation). Importantly, both $\theta$ and $\rho$ act indirectly through text observations rather than directly operating on the latent space. The fixed order simply determines how each value of $\theta$ or $\rho$ maps to a subset of semantic indices. For instance, omitting color in a caption corresponds to removing the associated latent variable—which index encodes this depends on the fixed ordering. Thus, adopting a graded lexicographic order ensures a reproducible indexing scheme without limiting model generality or expressiveness.
C.4 Proof of Corollary 4.2
We now proceed to prove Cor.˜4.2. For clarity, we restate the corollary below: See 4.2
@@
Given this structure, we directly apply the argument from the proof of Thm.˜4.1. The omission of variables in $\mathbb{I}^{2}_{\text{var}}$, together with the perturbation of variables in $\mathbb{I}^{1}_{\text{var}}$, ensures that only the invariant subset $\mathbb{I}_{\text{inv}}$ is both selected and unperturbed. Therefore, the invariant semantic components $\mathbf{s}_{\mathbb{I}_{\text{inv}}}$ are block-identified via smooth functions
@@
Appendix D Numerical Simulation Details
We provide additional details on the numerical simulations that are not fully covered in §˜5.1. Specifically:
  • In §˜D.1, we outline the experimental setup, including hyperparameters, model architectures, and the construction of downstream tasks.
  • In §˜D.2, we present additional experiments that further validate our theoretical findings.
  • In §˜D.3, we analyze downstream performance under various perturbation bias settings.
D.1 Detailed Experimental Setup
@@
Given the 10-dimensional semantic space, the total number of nonempty subsets is $2^{10}-1=1023$, defining the full range of selection bias configurations $\theta$. For tractability, we choose 10 representative selection subsets $\mathbb{I}_{\theta}$, detailed in Table˜2.
To isolate perturbation bias, we fix full selection ($\theta=1023$) and evaluate 10 representative perturbation subsets $\mathbb{I}_{\rho}$, as listed in Table˜3. Perturbations are introduced by independently modifying each semantic variable $s_{i}$ in $\mathbb{I}_{\rho}$ with probability 0.75 using additive Gaussian noise:
@@
where the joint noise covariance $\Sigma_{\epsilon}$ is sampled from $\mathcal{W}_{10}(I,10)$. Perturbations are applied solely to the text modality, aligning the proposed LVM and assumptions stated in §˜3.
@@
$\rho\mid\theta=1023$ 11 22 $12$ $57$ $177$ $387$ $639$ $849$ $969$ $1014$
$\mathbb{I}_{\rho}$ $\emptyset$ $\{1\}$ $[2]$ $[3]$ $[4]$ $[5]$ $[6]$ $[7]$ $[8]$ $[9]$
@@
We parameterize the generative functions $\mathbf{g}_{x}$ and $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ using randomly initialized 3-layer invertible MLPs, following prior work [13]. Invertibility is enforced by maintaining a condition number threshold of $1\mathrm{e}{-}3$ for each layer.
The encoding functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ are implemented as 7-layer MLPs and optimized using the Adam optimizer. For MMCL training, we use a batch size of $6144$, a learning rate of $1\mathrm{e}{-}4$, and train for 100,000100{,}000 iterations. The loss function is given by Eq.˜2, with Euclidean distance used as the similarity metric and a temperature parameter set to $1.0$. To ensure training stability, gradients are clipped using a maximum 2-norm of $2$.
@@
For both downstream tasks, we fix the pretrained encoders and evaluate the quality of the learned representations using a two-layer MLP as a probing model. We generate 20,48020{,}480 samples as the evaluation set for training the regressors and classifiers, along with an additional 20,48020{,}480 samples as the in-distribution test set. To assess OOD generalization, we generate another 20,48020{,}480 samples from the shifted latent space as the OOD test set.
The regressors are trained using Mean Squared Error (MSE) loss, and the classifiers use Cross-Entropy loss, both with a learning rate of $10^{-3}$ and trained for 10,00010{,}000 steps. We report classification performance using the average Matthews Correlation Coefficient (MCC). Importantly, in the OOD setting, we perform no adaptation or fine-tuning; the classifier is evaluated directly to test the generalization capability of the pretrained representations.
D.2 Additional Identification Results